# How does Labeling Error Impact Contrastive Learning? A Perspective from Data Dimensionality Reduction

### v1

@@ 3 Preliminaries
This section initiates our analysis with a comprehensive overview of contrastive learning. For any given samples $\bar{x}_{i},\tilde{x}_{i}\in\mathbb{R}^{d}$ and label $y_{i}\in\{1,...,K\}$, denote $\bar{\mathcal{D}}=\{\bar{x}_{i}\}_{i=1}^{n_{1}},\tilde{\mathcal{D}}=\{(\tilde{x}_{i},y_{i})\}_{i=1}^{n_{2}}$ as an original unlabeled dataset and a labeled dataset, respectively. Contrastive learning initializes and trains the model parameters in an unsupervised fashion, laying a foundational understanding of the underlying structure and patterns of data in $\bar{\mathcal{D}}$. The subsequent supervised fine-tuning stage further adapts the pre-trained model parameters to some specific downstream task on $\tilde{\mathcal{D}}$ to enhance model performance.
For the first learning stage, an encoder projector $f\in\mathcal{F}_{1}:\mathbb{R}^{d}\rightarrow\mathbb{S}^{k-1}$ is pre-trained to map a $d$-dimensional input vector to an embedding vector in a $k$-dimensional unit hypersphere, where the inequality $k<d$ generally holds which is validated by our experiments. This pre-training process is composed with several steps including data augmentation, contrastive representation, and loss calculation. Due to the unavailability of true label, contrastive pre-training constructs a surrogate task via some sample augmentation strategies. Then, this task is learned by minimizing the distance between similar samples and maximizing the distance between dissimilar samples in embedding space. For example, we first select an original sample $\bar{x}\in\bar{\mathcal{D}}$, and make data augmentations $t,t^{+}\in\mathcal{T}(\mathcal{T}=\{t|t:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\})$ to obtain two independent augmented samples $x=t(\bar{x}),x^{+}=t^{+}(\bar{x})$, respectively. Then, $M$ negative samples $x_{i}^{-},i=1,...,M$, are randomly augmented from other original samples in $\bar{\mathcal{D}}$. We assume $\bar{x},\tilde{x},x,x^{-}$ follow the same population distribution written as $p(x)$, which is the same as previous work like Arora et al. (2019); Wang et al. (2022). We use these augmented samples to calculate the most common contrastive loss, InfoNCE loss (van den Oord et al., 2018), defined as follows
@@
For the second learning stage, the parameters of the pre-trained encoder remain unaltered. We retrain a linear projection head $g\in\mathcal{F}_{2}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{K}$ with the weight $W\in\mathbb{R}^{k\times K}$ using the cross entropy (CE) loss to conduct downstream classification task. For a labeled sample $(\tilde{x},y)\in\tilde{\mathcal{D}}$, the CE loss and the mean CE loss  [^4] Mean classifier was first considered by Arora et al. (2019). They stated that the mean classifier could achieve comparable performance to learned weights. Note that we don’t use the mean classifier in our experiments. It is only available as an intermediate result (Theorem 4.2) in our theoretical analysis. are calculated via
$\mathcal{L}_{CE}(g_{f,W})=\mathbb{E}_{\tilde{x}}\left[-\log\frac{\exp\left(f(\tilde{x})^{\top}w_{y}\right)}{\sum_{i=1}^{K}\exp\left(f(\tilde{x})^{\top}w_{i}\right)}\right]$ (3)
@@
$\mathcal{L}_{CE}(g_{f,\mu})=\mathbb{E}_{\tilde{x}}\left[-\log\frac{\exp\left(f(\tilde{x})^{\top}\mu_{y}\right)}{\sum_{i=1}^{K}\exp\left(f(\tilde{x})^{\top}\mu_{i}\right)}\right],$ (4)
where the linear classifier predicts $g_{f,W}=\arg\max_{i\in[K]}\left(f^{\top}W\right)_{i},W=[w_{1},...,w_{K}]$, and the parameter of mean projection head is $\mu=[\mu_{1},...,\mu_{K}]$ whose element $\mu_{i}$ denotes the mean of the representations for the inputs with the label $i\in[K]$, i.e., $\mu_{i}=\mathbb{E}_{\{x|y=i\}}[f(x)]$. Define the downstream classification error as
$\mathcal{E}(f,W)=\underset{\tilde{x}\in\tilde{\mathcal{D}}}{\mathrm{Pr}}\left[g_{f,W}(\tilde{x})\neq y_{\tilde{x}}\right].$ (5)
@@
Assumption 3.1 (Labeling Error (HaoChen et al., 2021; Wang et al., 2024)).
For any $\bar{x}\in\bar{\mathcal{D}}$, its latent label $y_{\bar{x}}$, and its augmented sample $x\sim p(\cdot|\bar{x})$, we assume that the true label of $x$ is not consistent with $y_{\bar{x}}$ with the probability $\alpha\in(0,1)$. That is, $\mathbb{E}_{\bar{x}\in\bar{\mathcal{D}},x\sim p(\cdot|\bar{x})}\left[\mathbb{I}\left[y_{x}\neq y_{\bar{x}}\right]\right]=\alpha$.
@@ 4.1 Theoretical Impact of Labeling Error
Given a collection of augmentation strategies $\mathcal{T}$, we say that two original samples $\bar{x},\bar{x}^{\prime}\in\bar{\mathcal{D}}$ are $\mathcal{T}$-augmentation overlapped if they have overlapped views, i.e., $\exists t,t^{\prime}\in\mathcal{T}$ such that $t(\bar{x})=t^{\prime}(\bar{x}^{\prime})$.
@@
From the above results, it can be observed that there are multiple terms determining the upper and lower bounds of the mean classification risk $\mathcal{L}_{CE}(g_{f,\mu})$, i.e., 1) $\mathcal{L}_{InfoNCE}(f)$: the contrastive risk; 2) $V\left(f(x)\right)$: the intra-class variance of the representations of true positive augmented samples; 3) $V^{-}\left(f(x)\right)$: the variance of false positive augmented samples; 4) $V(f(x^{-}))=V(f|f\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})$: the intra-class variance of negative and true positive augmented samples; 5) $\mathcal{O}\left(M^{-\frac{1}{2}}\right)$: the order of the approximation error; 6) $\log\left(\frac{M}{K}\right)$: a constant depending on $M$ and $K$. In terms of the form of these bounds, our result is more refined than that of Theorem 4.2 in Wang et al. (2022) as evidenced by these terms $V\left(f(x)\right),V^{-}\left(f(x)\right)$, and $V(f(x^{-}))$. It should be noted that the assumption of our Theorem 4.2 is milder than the corresponding result in (Wang et al., 2022), which leads to some latent differences in these terms. We carry out the following analyses with respect to these terms.
@@ 4.2 Dimensionality Reduction as A New Perspective
This subsection investigates the impact of labeling error on contrastive learning through a new perspective of data dimensionality reduction. We take Singular Value Decomposition (SVD, Wall et al. (2003)) as an illustrative example. For ease of calculation, the randomized SVD (RSVD, Halko et al. (2011)) algorithm is employed in our experiments.
@@
In general, we use the truncated version of SVD which is expressed as $\hat{X}_{q}=U_{q}S_{q}V_{q}^{\top}$, where $U_{q}=[u_{1},...,u_{q}]\in\mathbb{R}^{m\times q},S_{q}=diag(s_{1},...,s_{q}),V_{q}=[v_{1},...,v_{q}]\in\mathbb{R}^{m^{\prime}\times q},q\in[m]$. The following lemma provided by Eckart & Young (1936) proved that $\hat{X}_{q}$ is the best least squares lower rank approximation with rank $q$ for a given matrix $X$.
@@
As stated by Kilmer et al. (2021), Lemma 4.4 implies that the majority of the informational content is captured by the dominant singular subspaces, i.e., the span of the singular vectors corresponding to the largest singular values. By default, we assume a positive correlation between the amount of information and its semantic relevance. It follows that the information content corresponding to the largest singular value $s_{1}$ should represent the most crucial information in $X$. For example, an image is intrinsically represented as a matrix $X$. The information associated with $s_{1}$ is most significant for distinguishing the true semantic of the image. Table 1 and Figure 3 empirically validate that the larger the singular value is, the more semantic-related information it encompasses. Figure 4 visually presents the differences between original samples and the samples with data dimensionality reduction.
@@
The truncated SVD with $q=q^{*}$ is defined as the optimal truncated SVD. As indicated by Assumption 4.5, we assume the trend where $\alpha$ initially decreases and then increases as the decrease of $q$. In the first stage ($q\geq q^{*}$), the decrease of $\alpha$ derives from the removal of semantically irrelevant information. Once $q<q^{*}$, semantic-related information begins to be discarded, giving rise to the increase of $\alpha$. Table 2 and Figure 5 (Appendix G) present empirical performances applying the truncated SVD with different values of $q$, aligning with our analysis above. Notably, the values of $q^{*}$ vary across different settings. Table 3 further validates the effectiveness of SVD across various augmentation strategies  [^6] Explanations of these augmentation strategies in Table 3 are provided in Appendix D.
@@ 4.3 Further Understanding of Labeling Error
Given an original dataset $\bar{\mathcal{D}}$ and an augmentation collection $\mathcal{T}$, there exist $n$ augmented samples that form the augmentation dataset
$\mathcal{D}_{aug}=\{x|x=t(\bar{x}),\bar{x}\in\bar{\mathcal{D}},t\in\mathcal{T}\}.$
An augmentation graph $\mathcal{G}$ is obtained by taking the $n$ augmented samples as the graph vertices and assuming there exists an edge between two vertices $x,x^{\prime}\in\mathcal{D}_{aug}$ (if they can be generated from a random original sample $\bar{x}\in\bar{\mathcal{D}}$).
According to spectral graph theory, we define $A\in\mathbb{R}^{n\times n}$ as the adjacency matrix of the augmentation graph $\mathcal{G}$. For two augmented samples $x,x^{\prime}\in\mathcal{D}_{aug}$, the element $A(x,x^{\prime})$ denotes the marginal probability of generating $x,x^{\prime}$ from a random original sample $\bar{x}\in\bar{\mathcal{D}}$. Formally, $A(x,x^{\prime})=\mathbb{E}_{\bar{x}\in\bar{\mathcal{D}}}\left[p(x|\bar{x})p(x^{\prime}|\bar{x})\right]$. The corresponding normalized graph Laplacian matrix is $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, where $D$ denotes a diagonal degree matrix with the diagonal element $D_{x,x}=\sum_{x^{\prime}\in\mathcal{D}_{aug}}A(x,x^{\prime})$. The eigenvalues of $L$ are denoted as $\{\lambda_{i}\}_{i=1}^{n}$, where $0=\lambda_{1}\leq...\leq\lambda_{n}\leq 2$.  [^7] When taking the truncated SVD with hyperparameter $q$ on the original dataset $\bar{D}$, the corresponding symbols are $\mathcal{G}_{q},A_{q},D_{q},L_{q}$, and $\lambda_{i,q},i\in[n]$.
@@
Let Assumption 3.1 hold. For the empirical optimal encoder $f^{*}$ taking the truncated SVD with hyper-parameter $q$ on the original dataset $\bar{\mathcal{D}}$, there exists a linear head $W$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k,q})$ such that
@@ Appendix A Notations
$\bar{\mathcal{D}}$ the original unlabeled dataset defined as $\{\bar{x}_{1},...,\bar{x}_{n_{1}}\},\bar{x}_{i}\in\mathbb{R}^{d}$
$\tilde{\mathcal{D}}$ the labeled dataset defined as $\{(\tilde{x}_{1},y_{1}),...,(\tilde{x}_{n_{2}},y_{n_{2}})\},\tilde{x}_{i}\in\mathbb{R}^{d},y_{i}\in[1,...,K]$
$p(\bar{x}|y),p(\tilde{x}|y)$ the distributions of the samples $\bar{x},\tilde{x}$ conditioned on the label $y$
@@
$\mathcal{E}(f,W)$ the downstream classification error defined as $\underset{\bar{x}\in\bar{\mathcal{D}}}{\mathrm{Pr}}\left[g_{f,W}(\bar{x})\neq y_{\bar{x}}\right]$
$\alpha$ the labeling error defined as $\mathbb{E}_{\bar{x}\in\bar{\mathcal{D}},x\sim p(\cdot|\bar{x})}\left[\mathbb{I}\left[y_{x}\neq y_{\bar{x}}\right]\right]$
@@
$\mathcal{L}_{spe}$ spectral contrastive loss

### v2

@@ 3 Preliminaries
This section initiates our analysis with a comprehensive overview of contrastive learning. We first denote $\bar{D}$ the finite but exponentially large set set of all unlabeled original sample $\bar{x}\in\mathbb{R}^{d}$, and denote $y:\bar{D}\rightarrow[K]$ the ground-truth (deterministic) labeling function, then $y_{\bar{x}}\in\{1,...,K\}$. Let $\mathcal{P}$ be the population distribution of the original samples. Contrastive learning initializes and trains the model parameters in an unsupervised fashion, laying a foundational understanding of the underlying structure and patterns of original data. The subsequent supervised fine-tuning stage further adapts the pre-trained model parameters to some specific downstream task on the test data drawing from $\mathcal{P}$ to enhance model performance.
For the first learning stage, an encoder projector $f\in\mathcal{F}_{1}:\mathbb{R}^{d}\rightarrow\mathbb{S}^{k-1}$ is pre-trained to map a $d$-dimensional input vector $\bar{x}$ to an embedding vector $z=f(x)$ in a $k$-dimensional unit hypersphere, where the inequality $k<d$ generally holds which is validated by our experiments. This pre-training process is composed with several steps including data augmentation, contrastive representation, and loss calculation. Due to the unavailability of true label, contrastive pre-training constructs a surrogate task via some sample augmentation strategies. Then, this task is learned by minimizing the distance between similar samples and maximizing the distance between dissimilar samples in embedding space. For example, we first select an original sample $\bar{x}$, and make data augmentations $t,t^{+}\in\mathcal{T}(\mathcal{T}=\{t|t:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}\})$ to obtain two independent augmented samples $x=t(\bar{x}),x^{+}=t^{+}(\bar{x})$, respectively. We use $p(\cdot|\bar{x})$ to denote the distribution of the augmented sample for $\bar{x}$. Then, $M$ negative samples $x_{i}^{-},i=1,...,M$, are randomly augmented from other original samples. We assume $\bar{x},x,x^{+},x^{-}$ follow the same population distribution, which is the same as previous work like Wang et al. (2022). We use these augmented samples to calculate the most common contrastive loss, InfoNCE loss (van den Oord et al., 2018), defined as follows
@@
For the second learning stage, the parameters of the pre-trained encoder remain unaltered. We retrain a linear projection head $g\in\mathcal{F}_{2}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{K}$ with the weight $W\in\mathbb{R}^{k\times K}$ using the cross entropy (CE) loss to conduct downstream classification task. For a labeled sample $x\sim\mathcal{P}$, the CE loss and the mean CE loss  [^4] Mean classifier was first considered by Arora et al. (2019). They stated that the mean classifier could achieve comparable performance to learned weights. Note that we don’t use the mean classifier in our experiments. It is only available as an intermediate result (Theorem 4.2) in our theoretical analysis. are calculated via
$\mathcal{L}_{CE}(g_{f,W})=\mathbb{E}_{x}\left[-\log\frac{\exp\left(f(x)^{\top}w_{y_{x}}\right)}{\sum_{i=1}^{K}\exp\left(f(x)^{\top}w_{i}\right)}\right]$ (3)
@@
$\mathcal{L}_{CE}(g_{f,\mu})=\mathbb{E}_{x}\left[-\log\frac{\exp\left(f(x)^{\top}\mu_{y_{x}}\right)}{\sum_{i=1}^{K}\exp\left(f(x)^{\top}\mu_{i}\right)}\right],$ (4)
where the linear classifier predicts $g_{f,W}=\arg\max_{i\in[K]}\left(f^{\top}W\right)_{i},W=[w_{1},...,w_{K}]$, and the parameter of mean projection head is $\mu=[\mu_{1},...,\mu_{K}]$ whose element $\mu_{i}$ denotes the mean of the representations for the inputs with the label $i\in[K]$, i.e., $\mu_{i}=\mathbb{E}_{\{x|y_{x}=i\}}[f(x)]$. Define the downstream classification error as
$\mathcal{E}(f,W)=\underset{x\sim\mathcal{P}}{\mathrm{Pr}}\left[g_{f,W}(x)\neq y_{x}\right].$ (5)
@@
Assumption 3.1 (Labeling Error (Wang et al., 2024)).
For any $\bar{x}\sim\mathcal{P}$, its latent label $y_{\bar{x}}$, and its augmented sample $x\sim p(\cdot|\bar{x})$, we assume that the true label of $x$ is not consistent with $y_{\bar{x}}$ with the probability $\alpha\in(0,1)$. That is, $\mathbb{E}_{\bar{x}\sim\mathcal{P},x\sim p(\cdot|\bar{x})}\left[\mathbb{I}\left[y_{x}\neq y_{\bar{x}}\right]\right]=\alpha$.
@@ 4.1 Theoretical Impact of Labeling Error
Given a collection of augmentation strategies $\mathcal{T}$, we say that two original samples $\bar{x},\bar{x}^{\prime}\sim\mathcal{P}$ are $\mathcal{T}$-augmentation overlapped if they have overlapped views, i.e., $\exists t,t^{\prime}\in\mathcal{T}$ such that $t(\bar{x})=t^{\prime}(\bar{x}^{\prime})$.
@@
From the above results, it can be observed that there are multiple terms determining the upper and lower bounds of the mean classification risk $\mathcal{L}_{CE}(g_{f,\mu})$, i.e., 1) $\mathcal{L}_{InfoNCE}(f)$: the contrastive risk; 2) $V\left(f(x)\right)$: the intra-class variance of the representations of true positive augmented samples; 3) $V^{-}\left(f(x)\right)$: the variance of false positive augmented samples; 4) $V(f(x^{-}))=V(z|z\in\{f(x^{+}),f(x^{-})\},y_{x^{+}}=y_{x})$: the intra-class variance of negative and true positive augmented samples; 5) $\mathcal{O}\left(M^{-\frac{1}{2}}\right)$: the order of the approximation error; 6) $\log\left(\frac{M}{K}\right)$: a constant depending on $M$ and $K$. In terms of the form of these bounds, our result is more refined than that of Theorem 4.2 in Wang et al. (2022) as evidenced by these terms $V\left(f(x)\right),V^{-}\left(f(x)\right)$, and $V(f(x^{-}))$. It should be noted that the assumption of our Theorem 4.2 is milder than the corresponding result in (Wang et al., 2022), which leads to some latent differences in these terms. We carry out the following analyses with respect to these terms.
@@ 4.2 Dimensionality Reduction as A New Perspective
This subsection investigates the impact of labeling error on contrastive learning through a new perspective of data dimensionality reduction. Specifically, we use data dimensionality reduction on individual original sample to obtain the corresponding compressed sample which is then augmented to pre-train the model. We take Singular Value Decomposition (SVD, Wall et al. (2003)) as an illustrative example. For ease of calculation, the randomized SVD (RSVD, Halko et al. (2011)) algorithm is employed in our experiments.
@@
In general, we use the truncated version of SVD which is expressed as $\hat{X}_{q}=U_{q}S_{q}V_{q}^{\top}$, where $U_{q}=[u_{1},...,u_{q}]\in\mathbb{R}^{m\times q},S_{q}=diag(s_{1},...,s_{q}),V_{q}=[v_{1},...,v_{q}]\in\mathbb{R}^{m^{\prime}\times q},q\in[m]$. The data distribution after conducting truncated SVD is written as $\mathcal{P}_{q}$, then the corresponding labeling error is $\alpha_{q}=\mathbb{E}_{\bar{x}\sim\mathcal{P}_{q},x\sim p(\cdot|\bar{x})}\left[\mathbb{I}[y_{x}\neq y_{\bar{x}}]\right]$. The following lemma provided by Eckart & Young (1936) proved that $\hat{X}_{q}$ is the best least squares lower rank approximation with rank $q$ for a given matrix $X$.
@@
As stated by Kilmer et al. (2021), Lemma 4.4 implies that the majority of the informational content is captured by the dominant singular subspaces, i.e., the span of the singular vectors corresponding to the largest singular values. By default, we assume a positive correlation between the amount of information and its semantic relevance. It follows that the information content corresponding to the largest singular value $s_{1}$ should represent the most crucial information in $X$. For example, an image is intrinsically represented as a matrix $X$. The information associated with $s_{1}$ is most significant for distinguishing the true semantic of the image. Table 1 and Figure 3 empirically validate that the larger the singular value is, the more semantic-related information it encompasses. Figure 4 visually presents the differences between original samples and the compressed samples.
@@
The truncated SVD with $q=q^{*}$ is defined as the optimal truncated SVD. As indicated by Assumption 4.5, we assume the trend where $\alpha$ initially decreases and then increases as the decrease of $q$. In the first stage ($q\geq q^{*}$), the decrease of $\alpha$ derives from the removal of semantically irrelevant information. Once $q<q^{*}$, semantic-related information begins to be discarded, giving rise to the increase of $\alpha$. Table 2 and Figure 5 (Appendix G) present empirical performances applying the truncated SVD with different values of $q$, aligning with our analysis. Notably, the values of $q^{*}$ vary across different settings. Table 3 further validates the effectiveness of SVD across various augmentation strategies  [^6] Explanations of these augmentation strategies in Table 3 are provided in Appendix D.
@@ 4.3 Further Understanding of Labeling Error
Given an augmentation collection $\mathcal{T}$, there exist $n$ augmented samples that form the augmentation dataset
$\mathcal{D}_{aug}=\{x|x=t(\bar{x}),\bar{x}\sim\mathcal{P},t\in\mathcal{T}\}.$
An augmentation graph $\mathcal{G}$ is obtained by taking the $n$ augmented samples as the graph vertices and assuming there exists an edge between two vertices $x,x^{\prime}\in\mathcal{D}_{aug}$ (if they can be generated from a random original sample $\bar{x}\sim\mathcal{P}$).
According to spectral graph theory, we define $A\in\mathbb{R}^{n\times n}$ as the adjacency matrix of the augmentation graph $\mathcal{G}$. For two augmented samples $x,x^{\prime}\in\mathcal{D}_{aug}$, the element $A(x,x^{\prime})$ denotes the marginal probability of generating $x,x^{\prime}$ from a random original sample $\bar{x}\sim\mathcal{P}$. Formally, $A(x,x^{\prime})=\mathbb{E}_{\bar{x}\sim\mathcal{P}}\left[p(x|\bar{x})p(x^{\prime}|\bar{x})\right]$. The corresponding normalized graph Laplacian matrix is $L=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$, where $D$ denotes a diagonal degree matrix with the diagonal element $D_{x,x}=\sum_{x^{\prime}\in\mathcal{D}_{aug}}A(x,x^{\prime})$. The eigenvalues of $L$ are denoted as $\{\lambda_{i}\}_{i=1}^{n}$, where $0=\lambda_{1}\leq...\leq\lambda_{n}\leq 2$.  [^7] When taking the truncated SVD with hyperparameter $q$ on the unlabeled original sample $\bar{x}$, the corresponding symbols are $\mathcal{G}_{q},A_{q},L_{q}$, and $\lambda_{i,q},i\in[n]$.
@@
Let Assumption 3.1 hold. For the empirical optimal encoder $f^{*}$ taking the truncated SVD with hyper-parameter $q$ on unlabeled original dataset, there exists a linear head $W^{*}\in\mathbb{R}^{k\times K}$ with norm $\|W^{*}\|_{F}\leq 1/(1-\lambda_{k,q})$ such that
@@ Appendix A Notations
$\bar{\mathcal{D}}$ the original unlabeled dataset drawing from the distribution $\mathcal{P}$
$p(x|\bar{x})$ the distribution of the augmented sample $x$ conditioned on the unlabeled original sample $\bar{x}$
@@
$\mathcal{E}(f,W)$ the downstream classification error defined as $\underset{x\sim\mathcal{P}}{\mathrm{Pr}}\left[g_{f,W}(x)\neq y_{x}\right]$
$\alpha$ the labeling error defined as $\mathbb{E}_{\bar{x}\sim\mathcal{P},x\sim p(\cdot|\bar{x})}\left[\mathbb{I}\left[y_{x}\neq y_{\bar{x}}\right]\right]$
@@
$\mathcal{L}_{spe}$ the spectral contrastive loss